OpenStudy (anonymous):

Please Assist Logs solve the following equation exactly for x: log(12-4x)-log(1-x)=1

OpenStudy (anonymous):

log a- log b= log a/b.....so log 12-4x/1-x =1 ==>12-4x/1-x=10 solve for x.... (:

OpenStudy (anonymous):

do I solve for x by cross multiplying ? @divu.mkr

OpenStudy (anonymous):

yep.....like 12-4x=10(1-x)

OpenStudy (anonymous):

would the answer be 1/3 ?

OpenStudy (anonymous):

you hav to solve.... :D

OpenStudy (anonymous):

 I sthis correct @divu.mkr ?